# Magnet: Augmenting Generative Decoders with Representation Learning and Infilling Capabilities

### v1

@@ 3.3 Approach Overview
Figure 4 provides an overview of Magnet. Starting with a training example $\mathbf{x}{}$, the process unfolds in two parallel streams – (1) One or more contiguous spans of $M$ tokens in $\mathbf{x}{}$ are marked as span tokens, while a fraction of the remaining tokens (context tokens) is masked to form $\mathbf{x}{}^{\text{m}}$. (2) $\mathbf{x}{}$ is augmented to get $\mathbf{x}{}^{+}$. The input sequences $\mathbf{x}{}$, $\mathbf{x}{}^{\text{m}}$ and $\mathbf{x}{}^{+}$ are processed by the base decoder model to produce hidden states $\mathbf{h}{}$, $\mathbf{h}{}^{\text{m}}$ and $\mathbf{h}{}^{+}$. From $\mathbf{h}{}^{\text{m}}$, a language modeling head generates $\mathbf{y}{}^{\text{m}}$, which is used to compute $\mathcal{L}_{\text{MNTP}}$ and $\mathcal{L}_{\text{MSG}}$. Parallelly, $\mathbf{h}{}$ and $\mathbf{h}{}^{+}$ are processed using a projection head to get e and $\textbf{e}^{+}$, which are used to compute $\mathcal{L}_{\text{SSCL}}$. The overall loss function is given as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{MNTP}}+\lambda_{2}\mathcal{L}_{\text{SSCL}}+\lambda_{3}\mathcal{L}_{\text{MSG}}$ (5)
For processing $\mathbf{x}{}$ to get $\mathbf{x}{}^{+}$, the decoder utilizes a bidirectional attention mask, illustrated in Figure 2(b). While processing $\mathbf{x}{}^{\text{m}}$, the decoder employs an attention mask similar to those depicted in Figures 2(c) and 2(d). In some cases, when all input tokens are marked as span tokens, the attention mask reduces to causal attention, as shown in Figure 2(a).
@@ 4 Experiments
In this section, we show that applying Magnet augments a decoder-only LLM with representation learning and infilling capabilities. All training details are mentioned in Appendix A. Additionally, we present ablation experiments demonstrating the benefits of training a bidirectional model with a causal objective in Appendix C.
4.1 Word-Level Tasks
We evaluate the token-level representations on three tasks – (1) chunking, (2) named entity recognition, and (3) part-of-speech tagging – using the CoNLL-2003 dataset (Sang & Meulder, 2003). After applying the training objectives proposed in Section 3.2, we train a linear classifier on top of the frozen representations obtained from the last hidden state of the model. The word-level embeddings are obtained by averaging the representations of the tokens that make up that word. Further, the representation of the token at position $i$ is given by the embedding at position $i-1$. More specifically, if $w_{1}w_{2}w_{3}$ denotes the original input sentence, $(x_{w_{1}1},x_{w_{1}2},x_{w_{2}1},x_{w_{2}2},x_{w_{2}3},x_{w_{3}1})$ denotes its tokenized form, and $(y_{w_{1}1},y_{w_{1}2},y_{w_{2}1},y_{w_{2}2},y_{w_{2}3},y_{w_{3}1})$ denotes the model’s output, the representation for the word $w_{2}$ is computed as $(y_{w_{1}2}+y_{w_{2}1}+y_{w_{2}2})/3$.
[…]
Model Chunking NER POS-Tagging
@@ 4.2 Sentence-Level Tasks
We compare the text encoding capabilities of Magnet with other recently proposed methods for transforming decoder models into text encoders, viz. LLM2Vec (BehnamGhader et al., 2024) and Echo Embeddings (Springer et al., 2024). Table 2 shows the results on Semantic Textual Similarity (STS) task and Table 3 shows the results on clustering tasks. As can be seen, Magnet outperforms other adaptation methods on STS and clustering tasks. As previously noted, the fact that Magnet surpasses LLM2Vec suggests the potential benefit of a unified training approach. Additionally, Magnet not only achieves better performance on text encoding tasks but also provides decoder models with more capabilities (e.g., infilling), which its competitors lack.
@@
Table 5: Human evaluation results on the infilling tasks. The score denotes the percentage of infillings that were considered contextually appropriate by human evaluators.
@@
Unidirectional LLaMA-2-7B 53.50
Zero-Shot Setup 5.50
Five-Shot Setup 54.50
Magnet 62.00
@@ 4.3 Infilling Task
To test infilling capabilities, we evaluate the perplexity (PPL) of LLaMA-2-7B and Magnet-adapted LLaMA-2-7B on the test set of ROC Stories (Mostafazadeh et al., 2016) and Wikitext-103 (Merity et al., 2016). For ROC Stories, we randomly mask out a sentence from each 5-sentence story, while for Wikitext-103, we mask up to three spans with lengths ranging from 8 to 32 tokens. Following Donahue et al. (2020), we compute PPL only for the tokens comprising the original masked out spans. The results are presented in Table 5, and they show that the base model (LLaMA-2-7B) exhibits significantly higher perplexity for the masked spans compared to Magnet, demonstrating that Magnet effectively augments the base model with text infilling capabilities. This improvement is attributed to Magnet’s ability to incorporate all the surrounding contextual information when infilling text, which increases the likelihood of correctly predicting the original masked content.
We also conducted experiments using zero-shot and few-shot learning to enable LLaMA-2-7B to incorporate all the surrounding context when infilling a missing span. We explored various prompting strategies and found that while a zero-shot setup did not yield sensible infillings, a five-shot setup with descriptive prompts resulted in more context-aware infillings (refer Appendix B for details). For a comprehensive analysis, we conducted a human evaluation to compare the quality of infillings generated by the base model, its zero-shot variant, its few-shot variant, and its Magnet adaptation. In this evaluation, we randomly sampled 100 stories from the ROC Stories dataset, masked out one of their middle sentences, and tasked the models with infilling the missing sentence. Two human annotators on Amazon Mechanical Turk then independently assessed whether each generated sentence was contextually appropriate and contributed to a coherent story. The results are presented in Table 5, and they show that the infillings generated by Magnet-adapted model are significantly more coherent than those generated by the variants of the base model. We show some qualitative examples in Table 10, demonstrating that Magnet successfully delivers relevant and contextually appropriate infillings.
Figure 5: Analyzing the repetition problem in text generation. Both LLM2Vec and Magnet are applied for 3400 iterations.
@@
Figure 6: LLM2Vec adaptation increases text repetition with more training, while no such trend is observed for Magnet.
4.4 Open-Ended Text Generation and Repetition Problem
The repetition problem in text generation refers to the issue when generative models repeatedly produce the same phrases or sentences. Prior studies have identified that this issue often results from biases in the training data, limitations in the model’s design, or standard likelihood training and inference (Holtzman et al., 2019; Welleck et al., 2019; Fu et al., 2020; Xu et al., 2022). In our study, we find that when generative decoder models are adapted into text encoders by enabling bidirectional attention (BehnamGhader et al., 2024; Li & Li, 2023; Li et al., 2023), the issue of repetition is significantly worsened. For example, Table 11 shows the texts generated (using greedy decoding) by the original LLaMA-2-7B and its adapted version for text encoding via LLM2Vec (BehnamGhader et al., 2024). We observe noticeable repetitions in the text generated by LLM2Vec-adapted-Llama, although the fine-tuning data (Wikitext-103) had almost no sentence-level repetitions (0.02%).
For quantitative detection of repetitions in texts generated by an LLM and its bidirectional adaptations, we compute Rep-Sen and Rep-4 (as done by prior works analyzing the repetition problem (Holtzman et al., 2019; Welleck et al., 2019; Xu et al., 2022)):
@@ 4.4 Open-Ended Text Generation and Repetition Problem
Specifically, we create a prefix-dataset from the test sets of Wikitext-103 and ROC Stories, consisting of 5-word and single-sentence prefixes, respectively. The model is then tasked with autoregressively generating text based on these prefixes. Table 6 shows the repetition metrics for LLaMA-2-7B and its adaptations using LLM2Vec and Magnet. As can be seen, Magnet makes the base model significantly less prone to repeating sentences. For instance, for Wikitext-103, LLM2Vec makes LLaMA-2-7B 36.5 times more likely to repeat sentences, while Magnet only makes it 2.7 times more likely. Further, as shown in Figure 6, the repetition problem exacerbates with additional iterations of LLM2Vec training, whereas no similar trend is observed with Magnet.
We conjecture that LLM2Vec is significantly more prone to generating repetitive text because it exclusively focuses on learning representations with bidirectional attention. This training approach perhaps makes the decoder model somewhat similar to bidirectional LMs like BERT, which are known to repeat words when used for text generation (Table 11). Magnet solves this issue by having autoregressive text generation as one of the training objectives.
5 Text Generation and Knowledge Retention
In Table 6, we evaluate the impact of Magnet on the knowledge acquired by the LLM during its pretraining, as measured by the Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) benchmark. We find that adapting the base model using the Wikitext-103 dataset results in a performance decline of 1.6% on average. This degradation can be attributed to the dataset’s encyclopedic nature, potentially introducing a knowledge bias that narrows the model’s generalization capabilities. To mitigate this limitations, we also adapt the model by fine-tuning on the SlimPajama dataset (Soboleva et al., 2023), which comprises texts from more diverse data sources like Commoncrawl, C4, GitHub, Books, ArXiv, Wikipedia, and StackExchange. With SlimPajama adaptation, the performance drop on MMLU is only 0.4% on average. Some categories (like ’Others’) show even slightly improved performance. This suggests that the adaptation method itself does not have a limiting effect on the model’s knowledge and generative ability.
Table 6: Evaluating the impact of Magnet on LLaMa-2-7B’s performance on the MMLU benchmark.
Model Dataset Humanities STEM Social Sciences Others Average
LLaMa-2-7B Unknown 42.9 36.4 51.2 52.2 45.3
Magnet Wikitext 41.1 33.5 49.5 52.1 43.7
Magnet SlimPajama 41.7 35.8 50.9 52.8 44.9
@@ 6 Conclusion
In this work, we presented Magnet, a method to transform causal LLMs into text encoders and infilling language models with bidirectional context-capturing ability. Through extensive experiments, we show that Magnet uniquely equips LLMs with abilities that are beyond the scope of traditional text encoders or decoders. Thus, Magnet shows the potential to unify text generation and text encoding within a single framework. Future research could explore scaling Magnet to multimodal settings.
@@ Appendix A Training Details
Magnet fine-tunes LLaMA-2-7B using LoRA (Hu et al., 2021) with $r=16$ and $\alpha=32$. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1e-8$, apply bfloat16 quantization, and use scaled-dot-product attention (SDPA). All experiments are performed on a single NVIDIA A100 GPU, with the Magnet adaptation of LLaMA-2-7B taking approximately 7 hours. The hyperparameters for the different training objectives are as follows:
MNTP.  We train for 4200 iterations using the Wikitext-103 train set (Merity et al., 2016) with a batch size of 32, a learning rate of 3e-5, and a max sequence length of 512. We select 20% of the tokens for masking – 80% of the selected tokens are replaced with a [MASK] token, 10% tokens are replaced with a random token from the model’s vocabulary, and 10% tokens are left unchanged. For LLaMA-2-7B, we use "_" as the mask token.
SSCL.  We train for 800 iterations with a batch size of 64, a learning rate of 3e-5, and a max sequence length of 128. To extract representations we use the prompt "Given the sentence, find its representation:" and extract the representations corresponding to the last token. The training data is created from Wikitext (Merity et al., 2016) by extracting lines longer than 20 words and paraphrasing them for the positive examples. We set $\tau=0.1$ in equation 3.
MSG.  Similar to MNTP, we train for 4200 iterations using the Wikitext-103 train set (Merity et al., 2016) with a batch size of 32, a learning rate of 3e-5, and a max sequence length of 512. A training example can have up to 2 missing spans with span length ranging from 4 to 128 tokens.
Overall Loss.  For the first 3400 iterations, we optimize the loss (equation 5) with $\lambda_{1}=1$, $\lambda_{2}=0$, and $\lambda_{3}=1$, and for the next 800 iterations $\lambda_{1}=1$, $\lambda_{2}=9$, and $\lambda_{3}=1$. We initially train with only MNTP and MSG because these objectives help the model learn to capture future context, a capability the base model lacks.
@@ B.2 Few-Shot Evaluation
Table 7: Ablation analysis of the proposed training objectives to assess the potential benefits or downsides of unified training.
@@ Appendix C Training Objective Ablation Analysis
Appendix D Comparing MTP and MNTP Objectives
We conducted an experiment to test whether we can use LoRA to adapt the base LLM for $l$-to-$l$ prediction (similar to BERT). The training curves for masked token prediction (MTP) and masked number token prediction (MNTP) are shown in Figure 7. As illustrated, with MTP, the loss converges, but the evaluation accuracy for masked token prediction decreases. This likely occurs because the base model is trained to predict the $(l+1)^{th}$ token at position $l$, and shifting to $l$-to-$l$ prediction introduces a significant distributional shift that the model may struggle to accommodate swiftly. Thus, overall, we find that MNTP is a more effective objective for converting a decoder-only LLM into a text encoder. Additionally, MNTP aligns well with the causal MSG objective and paves way for a unified text generator and encoder.
Appendix E Limitations
While Magnet better preserves the open-ended generation capability of the base LLM compared to other bidirectional adaptation methods, it still reduces generation quality. For instance, fine-tuning LLaMA-2-7B with Magnet increases the test set perplexity (PPL) on Wikitext-103 from 6.4 to 7.6. Although qualitative analysis shows no major artifacts in the generated text, the model’s confidence in predicting the next word seems to be somewhat diminished.
In the infilling task, we only focus on augmenting the base LLM with the ability to consider all the surrounding information to produce contextually coherent infilling. We have observed that the quality of infilling decreases when using the Magnet-adapted LLM to infill very long sequences in the middle of a text. We believe this observation is a result of the parameters set for the MSG objective, and it can be addressed by the following three ways: (1) Instead of attempting to infill a long segment of say 128 tokens at once, divide it into smaller chunks, such as infill 32 tokens at a time; (2) Adjusting the parameters of the MSG task; (3) Fine-tuning the model specifically for infilling tasks.
Since the pre-training data for LLaMA-2-7B is not publicly available, there is a potential risk of test set contamination. To this end, we compare test set performances with the original LLaMA-2-7B and other recent methods that use the same base model (LLM2Vec and Echo Embeddings). This comparative analysis helps ensure that our results are not unduly influenced by data contamination. We leave it to the future works to establish and test on benchmarks that are surely not a part of the pre-training data of the LLMs.

### v2

@@ 3.3 Approach Overview
Figure 4 provides an overview of Magnet. Starting with a training example $\mathbf{x}{}$, the process unfolds in two parallel streams – (1) One or more contiguous spans of $M$ tokens in $\mathbf{x}{}$ are marked as span tokens, while a fraction of the remaining tokens (context tokens) is masked to form $\mathbf{x}{}^{\text{m}}$. (2) $\mathbf{x}{}$ is augmented to get $\mathbf{x}{}^{+}$. The input sequences $\mathbf{x}{}$, $\mathbf{x}{}^{\text{m}}$ and $\mathbf{x}{}^{+}$ are processed by the base decoder model to produce hidden states $\mathbf{h}{}$, $\mathbf{h}{}^{\text{m}}$ and $\mathbf{h}{}^{+}$. From $\mathbf{h}{}^{\text{m}}$, a language modeling head generates $\mathbf{y}{}^{\text{m}}$, which is used to compute $\mathcal{L}_{\text{MNTP}}$ and $\mathcal{L}_{\text{MSG}}$. Parallelly, $\mathbf{h}{}$ and $\mathbf{h}{}^{+}$ are processed using a projection head to get e and $\textbf{e}^{+}$, which are used to compute $\mathcal{L}_{\text{SSCL}}$. The overall loss function is:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{MNTP}}+\lambda_{2}\mathcal{L}_{\text{SSCL}}+\lambda_{3}\mathcal{L}_{\text{MSG}}$
For processing $\mathbf{x}{}$ and $\mathbf{x}{}^{+}$, the decoder uses a bidirectional attention mask (as shown in Figure 2(b)). For processing $\mathbf{x}{}^{\text{m}}$, the decoder employs an attention mask similar to those depicted in Figures 2(c) and 2(d). In some cases, when all input tokens are marked as span tokens, the attention mask reduces to causal attention, as shown in Figure 2(a).
@@ 4 Experiments
In this section, we demonstrate that Magnet enhances a decoder-only LLM with representation learning and infilling capabilities while preserving its original generative abilities. Specifically, we show that LLMs adapted with Magnet outperform the base model and other adaptation methods on representation learning tasks (Sections 4.1 and 4.2). We then highlight how Magnet significantly improves the LLM’s ability to infill missing spans (Section 4.3). Finally, we analyze the effect of Magnet on the original generative and reasoning capabilities of the LLM (Section 4.5). All training details are mentioned in Appendix A. Additionally, we present ablation experiments demonstrating the benefits of training a bidirectional language model with a causal objective in Appendix C.
It is to be noted that our goal is not to achieve state-of-the-art results on a specific benchmark. Instead, we aim to enhance a pretrained LLM with additional capabilities while preserving its original performance. Therefore, our main baselines are the base LLM and other methods that augment the same LLM with specific capabilities.
Model Chunking NER POS-Tags
@@
4.1 Word-Level Tasks
We evaluate the token-level representations on three tasks – (1) chunking, (2) named entity recognition, and (3) part-of-speech tagging – using the CoNLL-2003 dataset (Sang and Meulder, 2003). After applying the training objectives proposed in Section 3.2, we train a linear classifier on top of the frozen representations obtained from the last hidden state of the model. The word-level embeddings are obtained by averaging the representations of the tokens that make up that word. Further, the representation of the token at position $i$ is given by the embedding at position $i-1$.
[…]
@@ 4.2 Sentence-Level Tasks
We compare the text encoding capabilities of Magnet with other recently proposed methods for transforming decoder models into text encoders, viz. LLM2Vec (BehnamGhader et al., 2024) and Echo Embeddings (Springer et al., 2024). Table 2 shows the results on Semantic Textual Similarity (STS) task and Table 3 shows the results on clustering tasks. As can be seen, Magnet outperforms other adaptation methods on STS and clustering tasks. As previously noted, the fact that Magnet surpasses LLM2Vec suggests the potential benefit of a unified training approach.
@@
Unidirectional Llama-2-7B 53.5
Zero-Shot Setup 5.5
Five-Shot Setup 54.5
Magnet 62.0
Table 5: Human evaluation for the infilling tasks. The score denotes the percentage of infillings that were considered contextually appropriate by human evaluators.
@@ 4.3 Infilling Task
To test infilling capabilities, we evaluate the perplexity (PPL) of Llama-2-7B and Magnet-adapted Llama-2-7B on the test set of ROC Stories (Mostafazadeh et al., 2016) and Wikitext-103 (Merity et al., 2016). For ROC Stories, we randomly mask out a sentence from each 5-sentence story, while for Wikitext-103, we mask up to three spans with lengths ranging from 8 to 32 tokens. Following Donahue et al. (2020), we compute PPL only for the tokens comprising the original masked out spans. The results are presented in Table 4, and they show that the base model (Llama-2-7B) exhibits significantly higher perplexity for the masked spans compared to Magnet, demonstrating that Magnet effectively augments the base model with text infilling capabilities.
We also conduct experiments using zero-shot and few-shot learning to enable Llama-2-7B to incorporate all the surrounding context when infilling a missing span. We explore various prompting strategies and found that while a zero-shot setup did not yield sensible infillings, a five-shot setup with descriptive prompts resulted in more context-aware infillings (refer to Appendix B for details). For a comprehensive analysis, we conducted a human evaluation to compare the quality of infillings generated by the base model, its zero-shot variant, its few-shot variant, and its Magnet adaptation. In this evaluation, we randomly sampled 100 stories from the ROC Stories dataset, masked out one of their middle sentences, and tasked the models with infilling the missing sentence. Two human annotators on Amazon Mechanical Turk (with at least a high school diploma) then independently assessed whether each generated sentence was contextually appropriate and contributed to a coherent story. The results are presented in Table 5, showing that the infillings generated by Magnet-adapted model are more coherent than those generated by the base model. We show some qualitative examples of infilling in Table 11.
@@
Table 6: Analyzing the repetition problem. Both LLM2Vec and Magnet are applied for 3400 iterations.
Figure 5: LLM2Vec increases text repetition with more training, while no such trend is observed for Magnet.
Model HellaSwag BBH ARC NQ MMLU
Easy Challenge Humanities STEM Social Science Other
Llama-2-7B 75.51 33.57 73.95 44.28 24.02 43.27 36.09 53.04 54.84
Magnet 75.08 32.22 74.33 44.52 24.22 42.25 36.63 52.64 52.40
Table 7: Evaluating the impact of Magnet on Llama-2-7B’s performance across benchmarks. The metrics are computed using the LM Evaluation Harness (Gao et al., 2024). Due to the undisclosed evaluation prompts for Llama, reproducing the exact baseline results is difficult. We adopt the same setup for both Llama-2-7B and Magnet.
4.4 Repetition Problem
The repetition problem in text generation refers to the issue when generative models repeatedly produce the same phrases or sentences. Prior studies have identified that this issue often results from biases in the training data, limitations in the model’s design, or standard likelihood training and inference (Holtzman et al., 2019; Welleck et al., 2019; Fu et al., 2020; Xu et al., 2022). In our study, we find that when generative decoder models are adapted into text encoders by enabling bidirectional attention (BehnamGhader et al., 2024; Li and Li, 2023; Li et al., 2023), the issue of repetition is significantly worsened. For example, Table 12 shows the texts generated (using greedy decoding) by the original Llama-2-7B and its LLM2Vec adaptation (BehnamGhader et al., 2024). We observe noticeable repetitions in the text generated by LLM2Vec-adapted-LLaMA, although the fine-tuning data (Wikitext-103) had almost no sentence-level repetitions (0.02%).
@@ 4.4 Repetition Problem
We conjecture that LLM2Vec is significantly more prone to generating repetitive text because it exclusively focuses on learning representations with bidirectional attention. This training approach perhaps makes the decoder model somewhat similar to bidirectional LMs like BERT, which are known to repeat words when used for text generation (Table 12). Magnet solves this issue by having autoregressive generation as an objective.
4.5 Knowledge and Reasoning Tasks
We assess the effect of Magnet on the knowledge and reasoning capabilities acquired by the LLM during pretraining. Specifically, we evaluate its performance on HellaSwag (0-shot) (Zellers et al., 2019), BBH (3-shot) (Suzgun et al., 2022), ARC (0-shot) (Clark et al., 2018), MMLU (5-shot) (Hendrycks et al., 2021), and NaturalQuestions (5-shot) Kwiatkowski et al. (2019), covering commonsense reasoning and world knowledge.
For this evaluation, we fine-tune the model on the SlimPajama dataset (Soboleva et al., 2023), which includes diverse text sources such as CommonCrawl, C4, GitHub, Books, ArXiv, Wikipedia, and StackExchange. This choice ensures that the fine-tuning data resembles Llama-2-7B’s original pretraining distribution (despite its exact composition being unknown). By doing so, we mitigate potential biases introduced by highly structured datasets like Wikitext, which could favor Wikipedia-derived tasks such as NaturalQuestions while disadvantaging commonsense reasoning benchmarks like HellaSwag.
As shown in Table 7, Magnet has minimal impact on the model’s knowledge and reasoning capabilities. The minor variations observed can be attributed to differences in dataset composition during fine-tuning. Furthermore, to ensure a comprehensive evaluation, we assess the performance of the model fine-tuned with SlimPajama on representation learning tasks and find the results consistent with the metrics reported in Tables 1 and 2 (where Wikitext-103 was used for fine-tuning). Specifically, when fine-tuned with SlimPajama, the model attains 92.00% on chunking, 98.30% on NER, 93.21% on POS-tagging, and an average of 79.33% on the STS tasks.
@@ 5 Conclusion
In this work, we presented Magnet, a method to transform causal LLMs into text encoders and infilling language models with bidirectional context-capturing ability. Through extensive experiments, we show that Magnet uniquely equips LLMs with abilities that are beyond the scope of traditional text encoders or decoders. Thus, Magnet shows the potential to unify text generation and text encoding within a single framework.
@@ Appendix A Training Details
Magnet fine-tunes Llama-2-7B using LoRA (Hu et al., 2021) with $r=16$ and $\alpha=32$. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1e-8$, apply bfloat16 quantization, and use scaled-dot-product attention (SDPA). All experiments are performed on a single NVIDIA A100 GPU, with the Magnet adaptation of Llama-2-7B taking approximately 7 hours. We discuss the training dataset and hyperparameters for the different objectives/tasks below.
Datasets.  For representation learning and infilling tasks, we train on Wikitext-103 (Merity et al., 2016) to ensure a fair comparison with our baselines in Sections 4.1 and 4.2. For knowledge and reasoning tasks (Section 4.5), we use SlimPajama (Soboleva et al., 2023) to mitigate biases from highly structured datasets like Wikitext, which could favor Wikipedia-derived tasks. The WikiText-103 dataset is released under the Creative Commons Attribution-ShareAlike license, and SlimPajama is available under the Apache 2.0 license, both permitting use in open-source research.
MNTP.  We train for 4200 iterations using a batch size of 32, a learning rate of 3e-5, and a max sequence length of 512. We select 20% of the tokens for masking – 80% of the selected tokens are replaced with a [MASK] token, 10% tokens are replaced with a random token from the model’s vocabulary, and 10% tokens are left unchanged. For Llama-2-7B, we use "_" as the mask token.
SSCL.  We train for 800 iterations using a batch size of 64, a learning rate of 3e-5, and a max sequence length of 128. To extract representations we use the prompt "Given the sentence, find its representation:" and extract the representations corresponding to the last token. The training data is created by extracting lines longer than 20 words and paraphrasing them for the positive examples. We set $\tau=0.1$ in equation 3.2.2.
MSG.  Similar to MNTP, we train for 4200 iterations using a batch size of 32, a learning rate of 3e-5, and a max sequence length of 512. A training example can have up to 2 missing spans, with span length ranging from 4 to 128 tokens.
Overall Loss.  For the first 3400 iterations, we optimize the loss (equation 3.3) with $\lambda_{1}=1$, $\lambda_{2}=0$, and $\lambda_{3}=1$, and for the next 800 iterations $\lambda_{1}=1$, $\lambda_{2}=9$, and $\lambda_{3}=1$. Initially, we train with only MNTP and MSG, as these objectives help the model learn to capture future context—a capability the base model lacks. However, this choice mainly contributes to faster training, as similar results are obtained when training with all objectives from the start.
@@ B.2 Few-Shot Evaluation
Table 8: Ablation analysis of the proposed training objectives to assess the potential benefits of unified training.
@@ Appendix C Training Objective Ablation Analysis
Appendix D Comparing MTP and MNTP Objectives
We conducted an experiment to test whether we can use LoRA to adapt the base LLM for $l$-to-$l$ prediction (similar to BERT). The training curves for masked token prediction (MTP) and masked number token prediction (MNTP) are shown in Figure 6. As illustrated, with MTP, the loss converges, but the evaluation accuracy for masked token prediction decreases. This likely occurs because the base model is trained to predict the $(l+1)^{th}$ token at position $l$, and shifting to $l$-to-$l$ prediction introduces a significant distributional shift that the model may struggle to accommodate swiftly. Thus, overall, we find that MNTP is a more effective objective for converting a decoder-only LLM into a text encoder. Additionally, MNTP aligns well with the causal MSG objective and paves the way for a unified text generator and encoder.
Appendix E Limitations
Given Llama-2-7B’s undisclosed pre-training data composition, there is a potential risk of test set contamination. We mitigate the undue influence of data contamination by benchmarking against the base model and other Llama-2-7B adaptations (LLM2Vec and Echo Embeddings). Future work should establish benchmarks guaranteed to be excluded from LLM pre-training data.
While Magnet preserves open-ended generation better than other bidirectional adaptation methods, it does impact the generation quality. For instance, fine-tuning Llama-2-7B with Magnet increases Wikitext-103 test-set perplexity from 6.4 to 7.6, indicating slightly reduced next-token prediction confidence despite maintaining artifact-free generation.
For infilling tasks, we focus solely on enabling the base LLM to leverage surrounding context for coherent text completion. The Magnet-adapted LLM shows reduced performance when infilling lengthy (more than 128 tokens) mid-text sequences. This limitation can be addressed by: chunking long infills into smaller segments, modifying MSG parameters, or task-specific fine-tuning for infilling.